An equal amount of money was spent on 4 similar lamps and 9 similar light bulbs . Each lamp coast $20 more than each light bulb.How much was spent together ?
I tried but I got remainders.

Answer by addingup(3677)

lamp = x
light bulb = y
x = y + 20
--------------------
4(y + 20) = 9y
4y + 80 = 9y
80 = 5y reverse the numbers to get the unknown on the left (looks better)
5y = 80
y = 16 the bulbs are $16 each
the lamps: 16+20 = $36 each
--------------------------------------
Check:
4(36) = 9(16)
144 = 144 Correct